OpenStudy (anonymous):

how many atoms are in 1.4 mol of PF(3)

OpenStudy (anonymous):

\[1.4 mole \times \frac{ 6.0223x10^{23} molecules PF _{3} }{mole } \times \frac{ 4 atoms }{ 1 molecule }\] =3.372x10^24 atoms